Eaton Electronic Company’s treasurer uses both the capital asset pricing model and the dividend valuation model to compute the cost of common equity (also referred to as the required rate of return for common equity).

Assume:

Rf = 8 %
Km = 13 %
β = 1.7
D1 = $ 0.90
P0 = $ 20
g = 9

a. Compute Ki (required rate of return on common equity based on the capital asset pricing model). (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.)

b. Compute Ke (required rate of return on common equity based on the dividend valuation model). (Do not round intermediate calculations. Input your answer as a percent rounded to 2 decimal places.)
Answer #1

a. Compute Ki (required rate of return on common equity based on the capital asset pricing model).

Answer: 16.50%

As per capital asset pricing model required rate of return on common equity is calculated using the following formula

Ki = Rf + (Km – Rf) * Beta

Data provided in the Question are

Rf = Risk Free Rate of Return = 8%

Km = Market Rate of Return = 13%

Beta = 1.7

Ki        = Rf + (Km – Rf) * Beta

            = 8% + (13% - 8%)*1.7

            =16.50%

b. Compute Ke (required rate of return on common equity based on the dividend valuation model).

Answer: 13.50%

Formula for calculating required rate of return on common equity (KE) based on the dividend valuation model is as follows

Ke = pi +g

Data provided in the Question are

D1 = Dividend in year 1 = $0.90

P0 = Price in year 0 = $20

g = Growth rate = 9% (i.e. 0.09)

Ke = (0.90/20) + .09

            = 13.50%

Note: in the question ''g'' is provided as 9, not as percentage so it is assumed as 9% (i.e. 0.90) and problem is solved accordingly.